# Task Preference Optimization: Improving Multimodal Large Language Models with Vision Task Alignment

### v1

@@ 4 Experiment
Our model is trained on a variety of visual task datasets and conversation datasets, as shown in Table 1. More training details are provided in the Appendix.
@@
We evaluate our given TPO on both general image/video understanding benchmarks (mainly in multiple choice form) and specific visual tasks (e.g. grounding, tracking, and so on). Specifically, we run our model along with other approaches on MVBench, VideoMME, NExT-GQA, MLVU, MMIU, and SEED-Bench2. MVBench [48] is designed to evaluate multi-modal fine-grained video understanding tasks for clips (lasting around 8 to 16 seconds), consisting of 20 video tasks relying heavily on temporal perception that cannot be addressed by single-frame analysis. Video-MME [22] is for evaluating MLLMs in both perception and reasoning across varying lengths of videos, annotated by humans. NExT-GQA [95] builds on NExT-QA by introducing timestamps that are crucial for understanding questions and determining answers. This requires MLLM to perform multi-step reasoning and emphasize a deeper understanding of both visual and textual content. MLVU [117] is constructed from a wide variety of long videos, with lengths ranging from 3 minutes to 2 hours, and includes nine distinct evaluation tasks. MMIU [60] is a comprehensive evaluation suite designed to assess LVLMs across a wide range of multi-image tasks. It encompasses 7 types of multi-image relationships, 52 tasks, 77K images, and 11K meticulously curated multiple-choice questions. SEED2-Bench2 [41] is a comprehensive benchmark for evaluating MLLMs, featuring 24K multiple-choice questions with human annotations. It spans 27 evaluation dimensions, assessing both text and image generation.
Concerning visual tasks, we test spatial grounding, moment retrieval, highlight detection, tracking, and referring segmentation, including 7 benchmarks and the results from several corresponding state-of-the-art expert models.
@@ 4.1 General Understanding Evaluation
TPO improves its baseline (VideoChat2) on several video understanding benchmarks with a notable margin. As shown in Table 2, VideoChat-TPO, using a 7B LLM and only 16 input frames, achieves a 66.8 average score on MVBench [48], increasing by 6.4% over VideoChat2 and exceeding the performance of ST-LLM [52] which uses 64 frames. Considering MVBench focuses on characterizing subtle temporal variations involving fine-grained action types, action order, moving direction, and so on, TPO’s seamless integration of detailed video understanding (like spatiotemporal grounding and tracking) into its optimized MLLM makes it effectively handle these video reasoning tasks.
On VideoMME [22], VideoChat-TPO outperforms comparable models, achieving a 9.3% improvement over VideoChat2 and demonstrating significant gains across short and medium-length videos, regardless of subtitle availability. On MLVU [117], VideoChat-TPO achieves a M-AVG score of 54.7, surpassing VideoChat2 by 10.2%. This confirms the effectiveness of TPO’s enhanced perceptions of long-form video evaluations. These results across three benchmarks demonstrate the significant advancements in multimodal video understanding achieved through TPO.
@@
Table 4 shows that VideoChat-TPO outperforms other models, achieving superior accuracy (Acc) and intersection over union (IoU) scores in NExT-GQA [95]. Its intersection over prediction (IoP) scores are comparable to those of LLoVi [110], which employs large, closed-source commercial models like GPT-4 [1]. The high Acc@IoP score not only reflects the enhanced capability of TPO-optimized VideoChat in effectively understanding and interpreting video content but also demonstrates its robustness in handling complex reasoning scenarios. Furthermore, the higher Acc@GQA score indicates that the model successfully integrates fine-grained temporal grounding with complex reasoning tasks, enabling it to accurately provide temporal clues necessary for inferring answers. The TPO training framework significantly enhances the reasoning capabilities of the baseline model, allowing it to overcome the limitations of traditional QA-trained models in identifying and localizing temporal elements within video data. This enhancement positions MLLM-TPO with a stronger competitive edge in the field of video reasoning.
@@
To explore TPO’s effect on multi-image understanding, we test it on the MMIU and SEED-Bench2, as shown in Table 14. On MMIU, VideoChat-TPO achieved an overall score of 40.2, a 5.2% improvement over VideoChat2. Besides temporal changes, MMIU evaluates models’ spatial sensing and semantic relations in scenes. On SEED-Bench2, across 27 performance indicators, VideoChat-TPO achieves 41.7% improvement on average performance compared to videochat2. VideoChat-TPO’s superior performance on multi-image understanding compared to LLaVA-Interleave and InternVL1.5-Chat demonstrates that TPO’s vision enhancements improve the MLLM’s spatial perception and image understanding.
@@ 4.2 Vision Task Evaluation
Moment Retrieval is to locate the target segments in a video based on a language query. Table 7 and Table 7 compare the zero-shot and fine-tuned moment retrieval performance of VideoChat-TPO against other expert models and MLLMs. In zero-shot settings on the Charades-STA dataset [24], VideoChat-TPO achieves a R@1 (IoU=0.5) score of 40.2, surpassing the previous state-of-the-art (SOTA) end-to-end trained MLLM, ChatVTG [66], and the expert model, UniVTG [50]. This demonstrates VideoChat-TPO’s ability to accurately locate video moments corresponding to given text queries, thereby enhancing the practical applicability of MLLMs.
@@
Highlight detection generates salient scores for emphasizing frames based on the input language query. We compare the fine-tuning highlight detection performance of VideoChat-TPO with that of other expert models and MLLMs, as shown in Table 7. VideoChat-TPO surpasses notably the previous state-of-the-art MLLM (such as TimeChat [72]) on both Charades-STA and QVHighlight. For expert models like QD-DETR and UniVTG, VideoChat-TPO beats them on Charades-STA non-trivially and achieves a comparable performance on QVHighlight. This also demonstrates the progress in extending MLLMs to broad temporal tasks.
@@
To verify the fine-grained localization ability of the model, we run the spatial grounding task which inputs textual descriptions into the model and gets the corresponding bounding box on RefCOCO [103]. As shown in Table 7, we compare VideoChat-TPO with pixel-to-sequence models, i.e., VisionLLM-H [84], pixel-to-embedding methods, i.e., NExT-Chat [109] and expert models, i.e., G-DINO [53]. We outperform the pixel-to-embedding methods by using only a simple task head and achieve comparable performance to pixel-to-sequence models fine-tuned on a large amount of spatial grounding data, as well as to specialized models.
@@
In the tracking task, the model receives the object coordinates from the first frame and outputs the coordinates for the remaining frames in the video. We evaluate the VideoChat-TPO on the mainstream tracking benchmarks LAOST [21] and GOT-10k [30], as shown in Table 10. In zero-shot testing, VideoChat-TPO ranks among the best of all MLLMs, even surpassing some fine-tuned expert models remarkably with around 10% increase on success rate, such as SiamFC[5] and ATOM [16]. With the TPO training method, the model is optimized from mask sequences that hard to be represented by language, allowing the MLLM to achieve strong motion characterization for multimodal generalization in highly dynamic and occluded scenes.
@@
The referring segmentation task requires the model to output the mask sequence corresponding to the specified prompt. This capability is not designed or learned in the original VideoChat2 or SAM2, but it can be activated in VideoChat-TPO. We compare the referring segmentation performance of VideoChat-TPO with other expert models and MLLMs in Tabel 10. The zero shot capability of VideoChat-TPO is comparable to the fine-tuning capabilities of other expert models, i.e. ReferFormer  [93], showing its notable open-world video segmentation performance. Through TPO training, MLLM effectively optimizes its understanding in both tracking and segmenting objects indicated by users. It further enables the model to excel in pixel-level tasks, offering perceptual advantages to it in practical applications like robotic control.
@@ 4.3 Ablation Studies
To evaluate the effectiveness of the TPO method on different MLLMs, we apply TPO to the LLaVA (LLaVA-OneVision [42]) in addition to VideoChat. In LLaVA, short-term fine-grained video understanding is a weakness. Despite the conversation data used in LLaVA-OneVision being very similar to, or even more extensive than, what we employ, there remains significant room for improvement in its performance on short-term fine-grained understanding tasks. As shown in Table 10, after optimization through TPO, LLaVA not only gains capabilities that it originally lacked but also achieves an 8.1% improvement on MVBench, demonstrating the universality of the TPO method.
@@
We compare the TPO method with the approach of textualizing task data and training it in an autoregressive manner, which is a straightforward and efficient method for MLLMs to learn specific tasks. We use the same data as the TPO method and convert the data related to the task head into conversation format. Due to the limitations of textual representation, much data must be input in a more discrete manner. For instance, in tracking supervision data, we follow Merlin [102] by converting masks into a sequence of spatial coordinates to serve as both the model’s input and output. As shown in Table 11, MLLM-TPO achieves a 2.1% performance improvement on MVBench [48] compared to MLLMs trained with textualized task data. This demonstrates that the task head in the TPO method enables more effective utilization of the original supervision signals compared to the next-token prediction approach, which inevitably incurs information loss when converting fine-grained tasks into textual annotations. The next-token prediction method cannot truly capture the nuanced information present in videos, while TPO allows the MLLM to learn perceptual information beyond conversational data.
@@
A stronger task head is expected to enhance the multimodal capabilities of TPO more effectively than a weaker one. To further explore the impact of the task head on model performance, we replaced the temporal head from CG-DETR [61] with a simple two-layer MLP. As shown in Table 11, the simple temporal head exhibits a significant performance decline compared to CG-DETR on the corresponding tasks. However, it still achieves a 1.1% improvement compared to the model without TPO. In contrast to the simpler head, the well-designed head can provide more accurate expert knowledge through its architecture and pre-trained task preference weights, enabling better utilization of the data.
@@
Enhancing model performance can be significantly achieved by incorporating additional data. In our approach, we expand the existing conversation data by integrating two reasoning datasets [7, 17] which offer valuable temporal information. With the addition of these datasets, we observe a significant enhancement in the capabilities of the temporal head in Table 11. Furthermore, as shown in Table 12, incorporating conversation data further improves performance across various tasks. Overall, increasing the amount of conversation and fine-grained task data leads to noticeable improvements in model performance, demonstrating the effectiveness of data scaling in training multimodal models.
Synergistic Gains from Co-training on Visual Tasks.
Table  12 presents the impact of incorporating different visual tasks into TPO on performance. This indicates that the collaborative learning of visual tasks facilitates the transfer of knowledge, ultimately leading to better performance across the board. Overall, these results confirm that co-training not only benefits individual tasks but also creates a synergistic effect that enhances overall capabilities in visual understanding.
@@ 5 Conclusions
This paper introduces Task Preference Optimization (TPO), which significantly enhances the overall multimodal performance of MLLMs by enhancing their precise understanding. TPO achieves this by integrating differentiable task preferences derived from fine-grained visual tasks. It introduces learnable task tokens that serve as bridges between multiple task-specific heads and the MLLM. Through the joint optimization of these task tokens, heads, and the MLLM, TPO leads to a substantial improvement in multimodal dialogue capabilities and performance across various visual tasks. Our results demonstrate the effectiveness of TPO in scaling MLLMs with task-specific data and seamlessly integrating them with existing expert vision models. We believe this work clarifies the prerequisites for fusing MLLMs with models and data from the pre-large model era, effectively bridging the gap between expert and generalist models, as well as between generation and understanding.
@@
Currently, TPO focuses exclusively on discriminative visual tasks, overlooking generative ones. Meanwhile, the framework is supervised by human annotations, which neglects potentially valuable unsupervised or self-supervised learning approaches such as contrastive learning [12, 26]. This limitation inherently restricts the scalability of TPO in terms of both task diversity and data requirements. While we demonstrate TPO’s potential for enhancing MLLMs through increased task-specific data, a comprehensive investigation into the broader impact of this scaling remains a crucial area for future research.
@@ 1 Detailed Results
The results are shown in Table 14. VideoChat-TPO shows a significant improvement over VideoChat2, achieving an overall score of 40.2 (a 5.2-point increase). Compared with VideoChat2, Our model has achieved clear improvements in Causality Reasoning (CR), Visually Grounded Reasoning (VGR), Multiple Image Captioning (MIC), Spot the Difference (STD), General Action Recognition (GAR), Temporal Localization (TL), Video Captioning (VidCap), Multiview Action Recognition (MAR), Image Captioning with Spatial Context (ICSC), and Egocentric Video Question Answering (EVQA), with scores of 73.0 (+26.5%), 69.5 (+15.2%), 83.0 (+19%), 92.5 (+61%), 88.0 (+15%), 94.5 (+13.0%), 73.4 (+35.8%), 48.5 (+12%) and 59.0 (14.5%), respectively. Among them, we suppose the improvement of TL capability comes from the optimization of our temporal head, and the improvement of VGR, STD, MAR and ICSC capabilities comes from the optimization of our region head and mask head. The enhancements observed in captioning, specifically in metrics such as MIC, IC, and VidCap, indicate an improvement of TPO to capture detailed visuals. Meanwhile, we find that the improvement in multi-image capabilities stems from enhanced instruction comprehension. Compared with video assessments, which primarily consist of multiple-choice questions, multi-image evaluations emphasize the accuracy of responses to specific questions. After optimization with TPO, the model has significantly improved its instruction following, leading to a higher success rate.

### v2

@@ 4 Experiment
Our model is trained on a variety of visual task datasets and conversation datasets, as shown in Table 1.
More training details are provided in the Appendix.
@@
We evaluate TPO on both general image/video understanding benchmarks (mainly in multiple choice form) and specific visual tasks (e.g. grounding, tracking, and so on). Specifically, we run our model along with other approaches on MVBench, VideoMME, NExT-GQA, MLVU, MMIU, and SEED-Bench2. MVBench [48] is designed to evaluate multi-modal fine-grained video understanding tasks for clips (lasting around 8 to 16 seconds), consisting of 20 video tasks relying heavily on temporal perception that can hardly be addressed by single-frame analysis. Video-MME [21] is for evaluating MLLMs in both perception and reasoning across varying lengths of videos, annotated by humans. NExT-GQA [90] builds on NExT-QA by introducing timestamps that are crucial for understanding questions and determining answers. This requires MLLM to perform multi-step reasoning and emphasize a deeper understanding of both visual and textual content. MLVU [112] is constructed from a wide variety of long videos, with lengths ranging from 3 minutes to 2 hours, and includes nine distinct evaluation tasks. MMIU [59] is a comprehensive evaluation suite designed to assess MLLMs across a wide range of multi-image tasks. It encompasses 7 types of multi-image relationships, 52 tasks, 77K images, and 11K meticulously curated multiple-choice questions. SEED2-Bench2 [40] is a MLLM benchmark, featuring 24K multiple-choice questions with human annotations. It spans 27 evaluation dimensions, assessing both text and image generation.
Concerning visual tasks, we test spatial grounding, moment retrieval, highlight detection, tracking, and referring segmentation, including 7 related benchmarks  [23, 38, 98, 20, 29, 17, 72] , and the results from several corresponding state-of-the-art expert models.
@@ 4.1 General Understanding Evaluation
TPO improves its baseline (VideoChat2) on several video understanding benchmarks with a notable margin. As shown in Table 2, VideoChat-TPO, using a 7B LLM and only 16 input frames, achieves a 66.8 average score on MVBench [48], increasing by 6.4% over VideoChat2 and exceeding the performance of ST-LLM [52] which uses 64 frames. Considering MVBench focuses on characterizing subtle temporal variations involving fine-grained action types, action order, moving direction, and so on, TPO’s seamless integration of detailed video understanding (like spatiotemporal grounding and tracking) into its optimized MLLM makes it effectively handle these fine-grained tasks.
On VideoMME [21], VideoChat-TPO outperforms the compared models, achieving a 9.3% improvement over VideoChat2 and demonstrating significant gains across short and medium-length videos, regardless of subtitle availability. On MLVU [112], VideoChat-TPO achieves a M-AVG score of 54.7, surpassing VideoChat2 by 10.2%. This confirms the effectiveness of TPO’s enhanced perceptions of long-form video evaluations. These results across three benchmarks demonstrate the notable advancements in multimodal video understanding achieved through TPO.
@@
Table 4 shows that VideoChat-TPO outperforms other models, achieving superior accuracy (Acc) and intersection over union (IoU) scores in NExT-GQA [90]. Its intersection over prediction (IoP) scores are comparable to those of LLoVi [105], which employs large, closed-source commercial models like GPT-4 [1]. The high Acc@IoP score not only reflects the enhanced capability of TPO-optimized VideoChat in effectively understanding and interpreting video content, but also demonstrates its robustness in handling complex reasoning scenarios. Furthermore, the higher Acc@GQA score indicates that the model successfully integrates fine-grained temporal grounding with complex reasoning tasks, enabling it to accurately provide temporal clues necessary for inferring answers. From this perspective, TPO training framework significantly allows MLLMs to overcome their limitations in identifying and localizing temporal elements within video data. This enhancement positions MLLM-TPO with a competitive edge in temporal-related reasoning.
@@
To explore TPO’s effect on multi-image understanding, we test it on the MMIU and SEED-Bench2, as shown in Table 4. On MMIU, VideoChat-TPO achieved an overall score of 40.2, a 5.2% improvement over VideoChat2. Besides temporal changes, MMIU evaluates models’ spatial sensing and semantic relations in scenes. On SEED-Bench2, across 27 performance indicators, VideoChat-TPO achieves 41.7% improvement on average performance compared to VideoChat2. VideoChat-TPO’s superior performance on multi-image understanding compared to LLaVA-Interleave and InternLM-XComposer demonstrates that TPO’s vision enhancements improve the MLLM’s spatial perception and image understanding.
@@ 4.2 Vision Task Evaluation
Moment Retrieval is to locate the target segments in a video based on queries. Table 7 and 7 compare the zero-shot and fine-tuned moment retrieval performance of VideoChat-TPO against other expert models and MLLMs. In zero-shot settings, VideoChat-TPO achieves a R@1 (IoU=0.5) score of 40.2 on the Charades-STA [23]. It surpasses the previous state-of-the-art (SOTA) MLLM specialized for temporal tasks, ChatVTG [65], and the expert model, UniVTG [50]. This demonstrates VideoChat-TPO’s ability to accurately locate video moments corresponding to given text queries, thereby enhancing the practical applicability of MLLMs.
@@
Highlight detection generates salient scores for emphasizing frames based on the input language query. We compare the fine-tuning highlight detection performance of VideoChat-TPO with that of other expert models and MLLMs, as shown in Table 7. VideoChat-TPO notably outperforms the previous state-of-the-art MLLM (such as TimeChat [70]) on both Charades-STA and QVHighlight. For expert models like QD-DETR and UniVTG, VideoChat-TPO beats them on Charades-STA non-trivially and achieves a comparable performance on QVHighlight. This also demonstrates the progress in extending MLLMs to broad temporal tasks.
@@
To verify the fine-grained localization ability of the model, we run the spatial grounding task which inputs textual descriptions into the model and gets the corresponding bounding box on RefCOCO [98]. As shown in Table 7, we compare VideoChat-TPO with pixel-to-sequence models, i.e., VisionLLM-H [82], pixel-to-embedding methods, i.e., NExT-Chat [104], and expert models, i.e., G-DINO [53]. VideoChat-TPO performs better than the pixel-to-embedding methods by using only a simple task head and achieves comparable performance to pixel-to-sequence models fine-tuned on a large amount of spatial grounding data, as well as to specialized models.
@@
In the tracking task, the model receives the object coordinates from the first frame and outputs the coordinates for the remaining frames in the video. We evaluate the VideoChat-TPO on the mainstream tracking benchmarks LAOST [20] and GOT-10k [29], as shown in Table 10. In zero-shot testing, VideoChat-TPO ranks best among all MLLMs, even surpassing some fine-tuned expert models notably with around 10% increase on success rate, such as SiamFC [5] and ATOM [15]. With the TPO training method, the model is optimized from mask sequences that hard to represent by words, allowing the MLLM to achieve strong motion characterization for multimodal generalization in highly dynamic and occluded scenes.
@@
The referring segmentation task requires the model to output the mask sequence corresponding to the specified prompt. This capability is newly activated in VideoChat-TPO, as it is not supported or designed in the original VideoChat2 or SAM2. We compare the referring segmentation performance of VideoChat-TPO with other expert models and MLLMs in Table 10. The zero shot capability of VideoChat-TPO is comparable to the fine-tuning capabilities of other expert models, i.e. ReferFormer  [88], showing its notable open-world video segmentation performance. Through TPO training, MLLM effectively optimizes its understanding in both tracking and segmenting objects indicated by users. It further enables the model to excel in pixel-level tasks, offering perceptual advantages to it in practical applications like robotic control.
@@ 4.3 Ablation Studies
To evaluate the effectiveness of the TPO method on different MLLMs, we apply TPO to LLaVA (LLaVA-OneVision [41]) in addition to VideoChat. In LLaVA, fine-grained video understanding is a weak spot. Despite the conversation data used in LLaVA-OneVision being very similar to, or even more extensive than, what we employ, there remains significant room for improvement in its performance on short-term fine-grained understanding tasks. As shown in Table 10, with TPO, LLaVA not only gains capabilities that it originally lacked but also achieves an 8.1% improvement on MVBench, demonstrating the universality of the TPO method.
@@
We compare the TPO method with the approach of textualizing task data and training it in an autoregressive manner, which is a straightforward and efficient method for MLLMs to learn specific tasks. We use the same data as the TPO method and convert the data related to the task head into conversation format. Due to the limitations of textual representation, much data must be input in a more discrete manner. For instance, in tracking supervision data, we follow Merlin [97] by converting masks into a sequence of spatial coordinates to serve as both the model’s input and output.
As shown in Table 11, MLLM-TPO achieves a 2.1% performance improvement on MVBench [48] compared to MLLMs trained with textualized task data.
This demonstrates that the task head in the TPO method enables more effective utilization of the original supervision signals compared to the next-token prediction approach, which inevitably incurs information loss when converting fine-grained tasks into textual annotations. The next-token prediction method cannot truly capture the nuanced information present in videos, while TPO allows the MLLM to learn perceptual information beyond conversational data.
@@
A stronger task head is expected to enhance the multimodal capabilities of TPO more effectively than a weaker one. To further explore the impact of the task head on model performance, we replaced the temporal head from CG-DETR [60] with a simple two-layer MLP.
As shown in Table 11, the simple temporal head exhibits a significant performance decline compared to CG-DETR on the corresponding tasks. However, it still achieves a 1.1% improvement compared to the model without TPO. In contrast to the simpler head, the well-designed head can provide more accurate expert knowledge through its architecture and pre-trained task preference weights, enabling better utilization of the data.
@@
Concerning foundation or large models, we expect to enhance their model performance by incorporating additional data at scale. In our approach, we expand the existing conversation data by integrating two reasoning datasets [6, 16] with temporal information. By incorporating these datasets, we observe a notable enhancement in the capabilities of the temporal head in Table 11. Furthermore, as shown in Table 12, adding conversation data further improves performance across various tasks. Overall, increasing the amount of conversation and fine-grained task data leads to noticeable improvements in model performance, demonstrating the effectiveness of data scaling in training multimodal models with TPO.
Synergistic Gains from Co-training.
Table 12 presents the impact of incorporating different visual tasks into TPO on performance.
This indicates that the collaborative learning of visual tasks facilitates the transfer of knowledge, ultimately leading to better performance across the board. Overall, these results confirm that co-training not only benefits individual tasks but also creates a synergistic effect that enhances overall capabilities in visual understanding.
@@ 5 Conclusions
This paper introduces Task Preference Optimization (TPO). It enhances the overall multimodal performance of MLLMs by enhancing their precise visual understanding. TPO achieves this by integrating differentiable task preferences derived from fine-grained visual tasks. It introduces learnable task tokens that serve as bridges between multiple task-specific heads and the MLLM. Through the joint optimization of these task tokens, heads, and the MLLM, TPO leads to a substantial improvement in multimodal dialogue capabilities and performance across various visual tasks. Our results demonstrate the effectiveness of TPO in scaling MLLMs with task-specific data and seamlessly integrating them with existing expert vision models. We believe this study clarifies the prerequisites for fusing MLLMs with models and data from the pre-large model era, effectively bridging the gap between expert and generalist models, as well as between generation and understanding.
@@
Currently, TPO focuses exclusively on discriminative visual tasks, overlooking generative ones. Meanwhile, the framework is supervised by human annotations, which neglects potentially valuable unsupervised or self-supervised learning approaches such as contrastive learning [12, 25]. This limitation inherently restricts the scalability of TPO in terms of both task diversity and requirements. While we demonstrate TPO’s potential for enhancing MLLMs through increased task-specific data, a comprehensive investigation into the broader impact of this scaling remains a crucial area for future research.
Acknowledgements
This work is partially supported by the National Key R&D Program of China (No. 2022ZD0160102), Jiangsu Frontier Technology R&D Program (No. BF2024076), and the Science and Technology Commission of Shanghai Municipality under Grant No. 23YF1461900.
@@ 1 Experiment Details
The results are shown in Table 14. VideoChat-TPO shows a significant improvement over VideoChat2, achieving an overall score of 40.2 (+5.2%). Compared with VideoChat2, Our model has achieved clear improvements in Causality Reasoning (CR), Visually Grounded Reasoning (VGR), Multiple Image Captioning (MIC), Spot the Difference (STD), General Action Recognition (GAR), Temporal Localization (TL), Video Captioning (VidCap), Multiview Action Recognition (MAR), Image Captioning with Spatial Context (ICSC), and Egocentric Video Question Answering (EVQA), with scores of 73.0 (+26.5%), 69.5 (+15.2%), 83.0 (+19%), 92.5 (+61%), 88.0 (+15%), 94.5 (+13.0%), 73.4 (+35.8%), 48.5 (+12%) and 59.0 (14.5%), respectively. Among them, we suppose the improvement of TL capability comes from the optimization of our temporal head, and the improvement of VGR, STD, MAR and ICSC capabilities comes from the optimization of our region head and mask head. The enhancements observed in captioning, specifically in metrics such as MIC, IC, and VidCap, indicate an improvement of TPO to capture detailed visuals. Meanwhile, we find that the improvement in multi-image capabilities stems from enhanced instruction comprehension. Compared with video assessments, which primarily consist of multiple-choice questions, multi-image evaluations emphasize the accuracy of responses to specific questions. After optimization with TPO, the model has significantly improved its instruction following, leading to a higher success rate.
@@
LLaVA-OV-TPO Performance on Video Benchmarks.
Model MVBench VideoMME PerceptionTest
LLaVa-OV 56.7 58.2 57.1
LLaVa-OV-TPO 64.8(+8.1) 61.3(+3.1) 64.0(+6.9)
Table 16: Perfermance of LLaVA-OV on Video Benchmarks.
According to Table 16, TPO method demonstrates performance improvements on LLaVA-OV [42] across multiple video benchmarks as it does in VideoChat [48] model. Since TPO uses extra visual cues to guide MLLM, LLaVA-OV-TPO achieves an average score of 64.8(+8.1%) on MVBench [48] and 64.0(+6.9%) on PerceptionTest [63]. The notable improvement clearly demonstrates the model’s greatly enhanced ability to perceive visual details. Moreover, LLaVA-OV-TPO achieves a 3.1% performance improvement on VideoMME [21] and shows that the model has also made progress in knowledge modeling and understanding long videos. These results suggest TPO method is effective across various models and is particularly beneficial for fine-grained perception tasks.